OpenStudy (anonymous):

Find the area of a rectangular object which has a length of 4sqrt2 inches and a width of 6sqrt6 inches. Explain, in complete sentences, how you arrived at the simplified answer and give the final solution in simplified radical form. What type of object in your home or school might this be?

OpenStudy (anonymous):

\[4\sqrt{2}\times 6\sqrt{6}\] \[=24\sqrt{12}\] \[=24\times 2\sqrt{3}\] \[=48\sqrt{3}\]

OpenStudy (anonymous):

ask your teacher how you are supposed to measure \[\sqrt{3}\] or \[\sqrt{6}\] or \[\sqrt{2}\] on a ruler

OpenStudy (anonymous):

i guess if you had a square with side 4, then the measure of the diagonal would be \[4\sqrt{2}\]

OpenStudy (anonymous):

Multiply what is outside of the sqrt then multiply what is under the sqrt. You'll get \[24\sqrt{12}\] then you simplify under the sqrt. If you break down 12 to prime numbers using a facter tree then you get 3*2*2. Since there is a group of 2 twos then you can take a 2 out and thats where the \[2\times24\sqrt{3}\] comes from.